# A Customized SAT-based Solver for Graph Coloring

### v1

@@ 3.6 Incremental bottom-up optimization
In particular, we make iterative calls to the SAT solver with increasing guesses $k$ of the chromatic number. SAT solvers do not allow the removal of clauses and thus constraints, but clauses can be deactivated through assumptions. The clause of Equation 8 that is used to prune subproblems can be extended with an activation literal $s_{k}$ as follows.
($\left(\bigvee_{\{u,v\}\in E(H)}e_{\operatorname{rep}(u)\operatorname{rep}(v)}\right)\lor s_{k}$ (9)
When looking for a $k$-coloring, the literal $s_{k}$ is assumed to be false so that Equation 9 will simplify to Equation 8 and can be used for pruning as normal. If we are not looking for a $k$-coloring anymore, we want to disable the clause and we do so by adding the assumption that the literal $s_{k}$ is true. The clause is now satisfied and will not cause any incorrect pruning.
@@ 4.1 Implementation Details of ZykovColor
Note that dominated vertices might occur again later during the Zykov recursion. Then, merging will be favored as a decision as described in Section 3.5.
@@
Our implementation uses the exact maximum clique algorithm CliSAT [^3] https://github.com/psanse/CliSAT [32] with a time limit of one second to compute a large initial clique, and an initial call to the Mycielskian lower bound procedure to compute a strong lower bound $lb$. These lower bounds are used in the iterative application of the reduction rules, after which the DSatur coloring heuristic is used for a quick upper bound $ub$ [7]. If this does not already solve the instance, the reduced graph is passed to the SAT solver.
@@
We determine the chromatic number using bottom-up search: We solve the $k$-coloring problem for $k=lb,lb+1,\ldots,ub-1$ until the first satisfiable problem is found. Performing bottom-up search has an advantage over top-down or binary search for our pruning-based algorithm since it allows us to always use the tightest possible upper bound $k$. The advantage of this is also confirmed experimentally in Section 4.3.
@@
For solving the $k$-coloring problem, we use the propagator as described in Section 3. In particular, it handles the transitivity constraints and lazily adds the needed clauses. It uses the clique and Mycielskian lower bounds to dynamically prune the search; both a greedy heuristic and local search are invoked to find large cliques. Further, positive pruning with cliques is enabled. Instead of a custom decision strategy we use the default strategy of CaDiCal, with the addition of deciding on dominated vertices first, if available. The SAT feature of assumptions is used to enable incremental bottom-up solving which allows to reuse information in the next decision problem. Finally, we use version 2.1.2 of CaDiCal as the underlying satisfiability solver. We refer to our algorithm in the configuration as described above as ZykovColor. The C++ code of the algorithm, including the propagator, data structures, and ported version of MNTS, is available at https://github.com/trewes/ZykovColor.
@@ 4.2 Test Setup
The DIMACS benchmark set [23] is the standard benchmark set for graph coloring. It comprises 137 real-world graphs, graphs with a certain structure, and random graphs. We use it as the main benchmark set.
@@
The first choice is the original implementation of the Zykov propagator gc-cdcl [^4] https://bitbucket.org/gkatsi/gc-cdcl/src/master/ [19], which uses the constraint programming solver MiniCSP [^5] https://bitbucket.org/gkatsi/minicsp. We also compare with two other Satisfiability algorithms: POP-S [^6] https://github.com/s6dafabe/popsatgcpbcp [11], using Kissat 3.1.1[5] to solve a partial order encoding of the problem, and CliColCom [^7] https://github.com/marijnheule/clicolcom [21], which alternates between computing larger cliques and colorings with fewer colors. To find improved colorings or prove optimality, they use either a local search solver or CaDiCal 1.5.2 on the assignment encoding.
Further, we also implemented the assignment encoding from Equations 1, 3 and 2. It is embedded in the same framework as ZykovColor, using the same SAT solver CaDiCal, initial lower and upper-bound heuristics, and initial reduction. We add the relatively weak symmetry-breaking constraints (10) and fix the color of the vertices in the clique $C$ found in preprocessing to exclude some symmetric solutions. This reduces the number of symmetries to $(k-|C|)!$.
@@
We refer to this algorithm as Assignment.
We also include the branch-and-price implementation exactcolors [^8] https://github.com/heldstephan/exactcolors [20], using Gurobi 12.0.0 as the underlying linear programming solver. One of the oldest paradigms for exact graph coloring is the branch-and-bound algorithm DSatur [7, 31]. We use the improved version [^9] https://bitbucket.org/gkatsi/gc-cdcl/src/master/sota/Segundo/DSATUR/ suggested by [31].
@@ 4.3 Ablation Study
We proposed several new features to obtain the ZykovColor algorithm and now want to analyze their individual contributions. Additionally, we want to evaluate the impact of search tree pruning. For that, we consider ZykovColor and the following configurations each having one feature changed or disabled, and two algorithms that use no pruning at all.
  • No MNTS: disabling the MNTS clique heuristic (described in Section 3.4)
  • Clique decision: Using the decision strategy of gc-cdcl (Section 4.4 of [19]) based on large cliques. It is motivated by DSatur’s branching heuristic.
  • No dominated: not using dominated vertices as decisions (Section 3.5), CaDiCal is making all decisions.
@@
  • Full Encoding: encoding with all clauses 5, 6 and 7 added, using the totalizer encoding [2] of Open-WBO [^10] https://github.com/sat-group/open-wbo [28] for the at-most-$k$ constraints.
  • Transitivity only: same as the Full Encoding, but instead of adding all transitivity clauses explicitly they are added and handled by the propagator.
Figure 3: Survival plot of ablation study for ZykovColor variants and Full Encoding/Transitivity only
Figure 3 provides a survival plot for the performance of the considered configurations on the DIMACS benchmark graphs. The $x$-axis shows the time limit in seconds, and the $y$-axis shows the number of solved instances within that time limit. Each marker represents an instance that was solved at the time of the $x$-coordinate. This allows one to compare the performance of different algorithms for any time limit of at most one hour and gives a quick overview of their performance profile.
We can see that pruning of the search tree during solving is the most essential feature of the propagator approach, as both the Full Encoding and the propagator that only handles the transitivity clauses solve far fewer instances. This shows the capabilities of using the IPASIR-UP interface provides. Instead of treating the SAT solver as a black box, we can significantly improve the performance by communicating information that is available to the user but not to the SAT solver. Between the Full Encoding and Transitivity only, we observe that even just handling the transitivity constraints with a propagator has an advantage over the full satisfiability encoding.
Considering the ZykovColor variants, we observe the configurations No MNTS, Top-down, and Clique-decisions to solve fewer instances than ZykovColor in one hour, with 90, 89, and 84 instances solved, respectively. The decision strategy appears to play a significant role. Despite the clique-based decision strategy being designed for the graph coloring problem, the general SAT solver strategy performs much better here. Regarding configurations No MNTS and Top-down, both lead to less pruning of the search space and the performance decreases noticeably.
The variants Non-incremental and No dominated perform similarly to ZykovColor and no clear winner can be deduced from the results as all algorithms solve 93 instances. ZykovColor has a higher median but lower average solving times compared with the other two and we, thus, choose this as the default configuration for our implementation.
@@ 4.4 Comparison with State-Of-The-Art Coloring Algorithms
DIMACS ZykovColor gc-cdcl POP-S Assignment CliColCom exactcolors DSatur
137 93 83 91 92 91 63 71
Table 1: Number of solved DIMACS instances within one hour. See LABEL:tab:dimacs-data in Appendix A for detailed results and individual runtimes.
Figure 4: Survival plot for different algorithms on the DIMACS graphs
Table 1 lists the number of solved instances out of the 137 graphs from the DIMACS benchmark set for each considered algorithm. A detailed table that lists the runtime for each solved instance can be found in LABEL:tab:dimacs-data. We visualize the results with a survival plot in Figure 4, which allows us to compare the performance for different time limits.
We can see that the SAT-based methods perform the best for any time limit. This is followed by the CSP/SAT hybrid gc-cdcl, the branch-and-bound method DSatur and the branch-and-price algorithm exactcolors.
ZykovColor, the algorithm based on the improvements presented in this paper, performs significantly better than gc-cdcl. This shows the positive impact of the new features presented in this work. Further, it solves two more instances than POP-S or CliColCom, the best-performing methods from the literature, and one more instance than the implementation of the assignment encoding. Only for small time limits are Assignment and CliColCom more successful than ZykovColor. In the end, all SAT methods perform similarly well, each solving between 91 and 93 instances.
The encoding plays an important role but differences in the preprocessing, the implementation, and the used solver also affect the performance. Notably, the Mycielskian bound used in the preprocessing of ZykovColor and Assignment can solve the usually difficult instance Myciel7, which POP-S and CliColCom cannot solve.
We briefly note that ZykovColor uses significantly less memory than Assignment and gc-cdcl on the DIMACS graphs; this is further discussed in Appendix B.
@@
Figure 5: Normalized algorithm performance by parameter for ER graphs
@@
We now compare the performance of different methods on the Erdős-Rényi graphs. We are particularly interested in density-dependent performances. To visualize this, we use a parallel coordinate plot in Figure 5. The set of instances is given on the $x$-axis, where $ER^{*}.p$ represents all Erdős-Rényi graphs of any size for the edge probability $p$. The $y$-axis represents the ratio of solved instances within the time limit of one hour.
[…]
@@
Looking at individual algorithm performance, we notice that DSatur and exactcolors, the worst two algorithms on the DIMACS graphs, are now performing much better. DSatur outperforms the SAT-based methods and gc-cdcl on all sets, and exactcolors become the best method for graphs of density $0.4$ and higher.
[…]
We conjecture that the lack of structure in Erdős-Rényi graphs is the reason the SAT-based methods are not competitive on most density classes. If the instance is very random, the learned conflict clauses are likely less useful. The reason for the superior performance of exactcolors on the dense instances could be the use of the fractional chromatic number as a lower bound. On dense graphs it becomes practically easy to compute and gives better lower bounds than a clique.
4.5 Additional Experiments on Open or Challenging Instances
[…]

### v2

@@ 3.6 Incremental Bottom-Up Optimization
In particular, we make iterative calls to the SAT solver with increasing guesses $k$ of the chromatic number. SAT solvers do not allow the removal of clauses and thus constraints, but clauses can be deactivated through assumptions. The clause of Equation˜8 that is used to prune subproblems can be extended with an activation literal $s_{k}$ as follows.
$\left(\bigvee_{\{u,v\}\in E(H)}e_{\operatorname{rep}(u)\operatorname{rep}(v)}\right)\lor s_{k}$ (9)
When looking for a $k$-coloring, the literal $s_{k}$ is assumed to be $\operatorname{False}$ so that Equation˜9 will simplify to Equation˜8 and can be used for pruning as normal. If we are not looking for a $k$-coloring anymore, we want to disable the clause and we do so by adding the unit clause $(s_{k})$. The clause of Equation˜9 is now satisfied and will not cause any incorrect pruning.
@@
3.7 Lower Bounds Using the Fractional Chromatic Number
The clique lower bound is not particularly strong, e.g., Mycielski graphs have clique number 2 and arbitrary high chromatic numbers. A stronger bound is the fractional chromatic number of a graph with a gap of $\mathcal{O}(\log n)$ [25].
The fractional chromatic number $\chi_{f}(G)$ is the value of the linear relaxation of the following stable set cover formulation, where $\mathcal{S}$ denotes the set of all stable sets in $G$.
$\chi(G)=$ $\min$ $\sum_{S\in\mathcal{S}}x_{S}$ (10)
s.t. $\sum_{S\in\mathcal{S}:v\in S}x_{S}\geq 1$ $\forall v\in V$ (CIP) (11)
$x_{S}\in\{0,1\}$ $\forall S\in\mathcal{S},$ (12)
Clearly, $\lceil\chi_{f}(G)\rceil\leq\chi(G)$ provides a lower bound.
As a consequence of [38], it is almost as hard to approximate as the chromatic number itself. However, in practice, it is often tractable using column generation [30, 20, 27].
Unlike the clique and Mycielskian bounds, this bound does not yield a specific subgraph that needs to be avoided. Thus, we simply add the trivial clause containing the negation of every decision in the current node to trigger the desired backtracking. This clause only prunes the current node and will not be useful in other parts of the search tree.
We use exactcolors [20] to compute the fractional chromatic number bound $\lceil\chi_{f}(G)\rceil$. Despite using a linear programming solver, it guarantees numerically safe lower bounds like all other methods used in our work. Computing the fractional chromatic number can be time-consuming. Usually, it is more efficient on dense graphs. Therefore, we only compute this bound after backtracking and if the density of the graph $G^{\prime}$ in the current Zykov node is greater or equal to $0.75$. The threshold was validated experimentally for the best performance profile on DIMACS benchmarks. On Erdős-Rényi graphs lower thresholds can be favorable as we will demonstrate in Section 4.4.
@@ 4.1 Implementation Details of ZykovColor
Note that dominated vertices might occur again later during the Zykov recursion. Then, merging will be favored as a decision as described in Section˜3.5.
@@
Our implementation uses the exact maximum clique algorithm CliSAT [^3] https://github.com/psanse/CliSAT [33] with a time limit of one second to compute a large initial clique, and an initial call to the Mycielskian lower bound procedure to compute a strong lower bound $lb$. The Dsatur coloring heuristic is used for a quick upper bound $ub$ [6]. If the instance is not yet solved, the lower bounds are used in the iterative application of the reduction rules, after which the lower and upper bound computations are called again if any vertices were removed. Additionally, exactcolors is called to compute the fractional chromatic number on the reduced graph with a time limit of 10 seconds. Should exactcolors not be able to report $\chi_{f}(G)$ within the 10 seconds, it is deemed too expensive to run inside the propagator, and the bounding technique described in Section˜3.7 is disabled. Finally, if this does not already solve the instance, the reduced graph is passed to the SAT solver.
@@
We determine the chromatic number using bottom-up search: We solve the $k$-coloring problem for $k=lb,lb+1,\ldots,ub-1$ until the first satisfiable problem is found. Performing bottom-up search has an advantage over top-down or binary search for our pruning-based algorithm since it allows us to always use the tightest possible upper bound $k$. The advantage of this is also confirmed experimentally in Section˜4.3.
@@
For solving the $k$-coloring problem, we use the propagator as described in Section˜3. In particular, it handles the transitivity constraints and lazily adds the necessary clauses. It uses the clique and Mycielskian lower bounds to dynamically prune the search; both a greedy heuristic and local search are invoked to find large cliques. Further, positive pruning using cliques as subgraph is enabled, see Lemma˜3.1. The fractional chromatic number is computed as lower bound in certain nodes, see Section˜3.7. Instead of a custom decision strategy, we use the default strategy of CaDiCaL, with the addition of deciding on dominated vertices first, if available. The SAT feature of assumptions is used to enable incremental bottom-up solving, which allows the solver to reuse information in the next decision problem. Finally, we use version 2.1.2 of CaDiCaL as the underlying satisfiability solver. We refer to our algorithm in the configuration as described above as ZykovColor (short name ZC).
@@ 4.2 Test Setup
The DIMACS benchmark set [23] is the standard benchmark set for graph coloring. It comprises 137 instances, which include real-world graphs, graphs with a certain structure, and random graphs. We use it as the main benchmark set.
@@
The first choice is the original implementation of the Zykov propagator gc-cdcl [^4] https://bitbucket.org/gkatsi/gc-cdcl/src/master/ [19], which uses the constraint programming solver MiniCSP [^5] https://bitbucket.org/gkatsi/minicsp. We also compare with two other satisfiability algorithms: POP-S [^6] https://github.com/s6dafabe/popsatgcpbcp [11], using Kissat 3.1.1777Replacing the solver with version 4 of Kissat or Version 2.1.2 of CaDiCaL resulted in fewer solved instances[5] to solve a partial order encoding of the problem, and CliColCom [^8] https://github.com/marijnheule/clicolcom (short name CCC) [21], which alternates between computing larger cliques and colorings with fewer colors. To find improved colorings or prove optimality, they use either a local search solver or CaDiCaL 1.5.2999Replacing the solver with version 2.1.2 of CaDiCaL resulted in errors for some instances and ultimately solved fewer instances on the assignment encoding.
Further, we also implemented the assignment encoding from (1)–(2). It is embedded in the same framework as ZykovColor, using the same SAT solver CaDiCaL, initial lower and upper-bound heuristics, initial reductions, and bottom-up search strategy. We add the relatively weak symmetry-breaking constraints (13) and fix the color of the vertices in the clique $C$ found in preprocessing to exclude some symmetric solutions. This reduces the number of symmetries to (k|C|)!(k-|C|)!.
@@
We refer to this algorithm as Assignment (Ass).
We also include the branch-and-price implementation exactcolors [^10] https://github.com/heldstephan/exactcolors (EC) [20], using Gurobi 12.0.0 as the underlying linear programming solver. Note that inside ZykovColor, we use exactcolors without branching, i.e., only to compute the fractional chromatic number. In the comparison here, it uses branching to compute the chromatic number. One of the oldest paradigms for exact graph coloring is the branch-and-bound algorithm DSatur [6, 32]. We use the improved version [^11] https://bitbucket.org/gkatsi/gc-cdcl/src/master/sota/Segundo/DSATUR/ suggested by [32].
@@ 4.3 Ablation Study
We proposed several new features to obtain the ZykovColor algorithm and now want to analyze their individual contributions. Additionally, we want to evaluate the impact of the search tree pruning. To this end, we consider ZykovColor and the following configurations each having one feature changed or disabled, and the last two variants that use no pruning at all.
  • No MNTS: disabling the MNTS clique heuristic (described in Section˜3.4)
  • Clique decision: Instead of the decision strategy described in Section˜3.5, use the strategy of gc-cdcl (Subsection 4.4 of [19]) based on large cliques. It is motivated by DSatur’s branching heuristic.
  • No dominated: not using dominated vertices as decisions (Section˜3.5), CaDiCaL is making all decisions.
@@
  • No-Fractional: disable the computation of the fractional chromatic number using exactcolors (Section˜3.7)
  • Full Encoding: encoding with all clauses (5)-(7) added, using the totalizer encoding [2] of Open-WBO [^12] https://github.com/sat-group/open-wbo [29] for the at-most-$k$ constraints. No lower bound pruning is applied.
  • Transitivity only: same as the Full Encoding, but instead of adding all transitivity clauses explicitly, they are added and handled by the propagator.
Figure 3: Survival plot of ablation study for ZykovColor variants and Full Encoding/Transitivity only.
Figure˜3 provides a survival plot for the performance of the considered configurations on the DIMACS benchmark graphs. The $x$-axis shows the time limit in seconds, and the $y$-axis shows the number of solved instances within that time limit. Each marker represents an instance that was solved at the time of the $x$-coordinate. This allows one to compare the performance of different algorithms for any time limit of at most one hour and gives a quick overview of their performance profile.
We can see that pruning of the search tree during solving is the most essential feature of the propagator approach, as both the Full Encoding and the propagator that only handles the transitivity clauses solve far fewer instances. This shows the capabilities of employing the IPASIR-UP interface. Instead of treating the SAT solver as a black box, we can significantly improve the performance by communicating information that is available to the user but not to the SAT solver. Between the Full Encoding and Transitivity only, we observe that even just handling the transitivity constraints with a propagator has an advantage over the full satisfiability encoding.
Considering the ZykovColor variants, we observe the configurations Top-down and Clique-decisions to solve fewer instances than ZykovColor in one hour, with 89 and 84 instances solved, respectively. The decision strategy appears to play a significant role; despite the clique-based decision strategy being designed for the graph coloring problem, the general SAT solver strategy performs much better here. Regarding Top-down, it leads to less pruning of the search space since a larger lower bound is required for a conflict, and the performance decreases noticeably.
The variants No Fractional, No MNTS, and No dominated perform similarly but slightly worse than ZykovColor and Non-incremental, solving 93 instances overall. The latter two configurations perform very similarly as well, and both solve 94 instances (the last instance is solved at the same time). ZykovColor shows slightly better performance for lower time limits and manages to solve some instances faster than Non-incremental. Thus, we choose ZykovColor as the default configuration for our implementation.
@@ 4.4 Comparison with State-Of-The-Art Coloring Algorithms
Table˜1 lists the number of solved instances out of the 137 graphs from the DIMACS benchmark set for each considered algorithm. A detailed table that lists the running time for each solved instance can be found in Table˜2 in the appendix. We visualize the results with a survival plot in Figure˜4, which allows us to compare the performance for different time limits.
We can see that the SAT-based methods perform the best for any time limit, followed by the CSP/SAT hybrid gc-cdcl, the branch-and-bound method DSatur and the branch-and-price algorithm exactcolors.
ZykovColor, the algorithm based on the improvements presented in this paper, performs significantly better than gc-cdcl. This shows the positive impact of the new features presented in this work. Further, it solves three more instances than POP-S or CliColCom, the best-performing methods from the literature, and two more instances than the implementation of the assignment encoding. Assignment and CliColCom are more successful than ZykovColor only for small time limits. ZykovColor outperforms CliColCom for time limits from 400 seconds and Assignment for time limits from 300 seconds. Ultimately, all SAT methods perform similarly well, each solving between 91 and 94 instances.
The encoding plays an important role, but differences in the preprocessing, the implementation, and the used solver also affect the performance. Notably, the Mycielskian bound used in the preprocessing of ZykovColor and Assignment can solve the difficult instance Myciel7, which POP-S and CliColCom cannot solve.
Figure 4: Survival plot for different algorithms on the DIMACS graphs.
Table 1: Number of solved DIMACS instances within one hour.
DIMACS ZC gc-cdcl POP-S Ass CCC EC DSatur
137 94 83 91 92 91 63 71
@@
We now compare the performance of different methods on the Erdős-Rényi graphs. We are particularly interested in density-dependent performances. To visualize this, we use a parallel coordinate plot in Figure˜5. The set of instances is given on the $x$-axis, where $ER^{*}.p$ represents all Erdős-Rényi graphs of any size for the edge probability $p$. The $y$-axis represents the ratio of solved instances within the time limit of one hour.
We can see that all algorithms except for exactcolors exhibit similar performance curves depending on the density parameter $p$. Initially, the performance decreases with increasing density until $ER^{*}.7$, after which the performance again increases for the very dense instances $ER^{*}.9$. Exactcolors follows a different trend. The performance is roughly constant on graphs $ER^{*}.2$ to $ER^{*}.5$. Then, it increases for $ER^{*}.7$ and $ER^{*}.9$. The sparsest instances $ER^{*}.05$ and $ER^{*}.1$ are easy for all algorithms.
@@
Looking at individual algorithm performance, we notice that DSatur and exactcolors, the worst two algorithms on the DIMACS graphs, are now performing much better. DSatur outperforms the SAT-based methods and gc-cdcl on all sets, and exactcolors becomes the best method for graphs of density $0.4$ and higher.
Assignment, POP-S, and ZykovColor perform similarly. Assignment is slightly better, and ZykovColor is slightly behind the two other methods. However, ZykovColor shows an advantage on very dense graphs. For the $ER^{*}.9$ graphs, the fractional chromatic number is used as bound after every backtrack (as the density of the graphs is always above the threshold of 0.75), which notably increases the performance compared to Assignment and POP-S.
CliColCom performs significantly differently compared to on the DIMACS graphs. It only shows good performance for the sparsest graphs up to $ER^{*}.2$.
As exactcolors performs very well on the Erdős-Rényi graphs, we attempt to leverage its strength by computing the fractional chromatic number inside ZykovColor more frequently. We do so by reducing the density threshold from $0.75$ to $0.4$, as exactcolors becomes the best algorithm at $ER^{*}.4$. This configuration is labeled ZykovColor-p4 in Figure˜5 and we can see that this considerably improves the ratio of solved instances for $ER^{*}.4$ to $ER^{*}.7$.
Figure 5: Normalized algorithm performance by parameter for ER graphs.
Impact of the Threshold on Erdős-Rényi Graphs
Using the fractional chromatic number lower bound with a smaller density threshold makes ZykovColor more competitive on Erdős-Rényi graphs than other SAT approaches. It is unclear how one would leverage this bound in the assignment or other related SAT formulations.
We did not see a benefit in using a lower density threshold on the DIMACS instances. The running time spent on computing fractional lower bounds rather lowers the performance profile.
However, on Erdős-Rényi graphs exactcolors remains the leading algorithm. We conjecture that one reason is the lack of structure in the random graphs. Here, the learned conflict clauses are likely less useful, and a strong bound such as the fractional chromatic number might be more decisive.
Another aspect could be the branching strategy. In exactcolors, the fractional covering solution is used to make branching decisions, which shows good results for the random Erdős-Rényi graphs. Modern SAT solvers use activity scores of the variables for their branching decision, which are influenced by the learned clauses. Again, if the instance is random and the learned clauses are less meaningful, the decisions made in the SAT solver might not be as good as those made by exactcolors. On the instance DSJC125.9, exactcolors visits 114 Zykov nodes, while ZykovColor visits 7400 nodes. Using the fractional chromatic number bound in every Zykov tree node and not only after backtracking would reduce this to 1000 nodes.
Memory Usage
On the DIMACS graphs, ZykovColor, Assignment, and gc-cdcl used less than 4.5, 7.5, and 8.5 GB, respectively, except for the instance qg.order100, where the consumption was 25, 25, and 44 GB. The average consumption was 385, 685, and 825 MB.
It is surprising that Assignment uses more memory than ZykovColor, despite the compactness of the encoding. The Zykov-based encoding needs $O(n^{2})$ variables, one for each non-edge of the graph. This could become significant for large sparse graphs. Despite many such graphs in the benchmark set, ZykovColor requires noticeably less memory. This indicates a more effective traversal of the search tree and potentially fewer or shorter added conflict clauses.
For the small sizes of the Erdős-Rényi graphs, the algorithms ZykovColor, Assignment, and gc-cdcl all use about 60 MB of memory on average.